At 99゚C and 748 Torres, a sample of a volitile liquid is vaporized completely in a 250 mL flask. The condensed vapor weighs 1.097 g calculate the molar mass of the gas
Answer #1

Given:

P = 748.0 torr

= (748.0/760) atm

= 0.9842 atm

V = 250.0 mL

= (250.0/1000) L

= 0.25 L

T = 99.0 oC

= (99.0+273) K

= 372 K

find number of moles using:

P * V = n*R*T

0.9842 atm * 0.25 L = n * 0.08206 atm.L/mol.K * 372 K

n = 8.06*10^-3 mol

mass(solute)= 1.097 g

use:

number of mol = mass / molar mass

8.06*10^-3 mol = (1.097 g)/molar mass

molar mass = 1.361*10^2 g/mol

Answer: 1.36*10^2 g/mol